OpenStudy (anonymous):

You Answer It Ill Fan Every Last one of You But i Will Medal Who Helps Me.. 2/5 t = 12/15

OpenStudy (anonymous):

Well if 2/5 t = 12/15.. 12/15 = 4/5 so t = (4/5)/(2/5) so t = 2